OpenStudy (anonymous):

Hi there. Brenda deposited some money in a bank. The bank paid her a simple interest every year. After t years, the amount of money f(t), in dollars, that the bank paid Brenda was $2,000 greater than 10 times the number of years. Which statement best describes the steps to sketch a graph showing the relationship between f(t) and t? Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 2000 and y-intercept = 10. Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 10 and y-intercept = 2000. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 10 and y-intercept = 200. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 2000 and y-intercept = 200. --------------------------------------------------------------------------------

OpenStudy (anonymous):

I got another question after this, too.

OpenStudy (anonymous):

@Vocaloid @mathstudent55

OpenStudy (anonymous):

Oh I might as well just post it already... Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (xapproachesinfinity):

10 times the number of years can be written as 10t 2000 greater than 10 times the number of years t Is written 10t+2000

OpenStudy (xapproachesinfinity):

So f(t) = 10t+2000

OpenStudy (anonymous):

So it would be... A?

OpenStudy (xapproachesinfinity):

Yes

OpenStudy (anonymous):

Thank you, can you help with the second question?

OpenStudy (xapproachesinfinity):

Oh b not a my dear

OpenStudy (xapproachesinfinity):

I gotta go sorry

OpenStudy (anonymous):

Oh okay, thank you anyway

OpenStudy (xapproachesinfinity):

Np